The production department for a company requires 7200kg of raw material for manufacturing
particular item per year. It has been estimated that the cost of placing an order is A 72 and the
cost of carrying inventory is 25 percent of the investment in the inventories. The price of raw
material is A 20 per kg. Find the economic order quantity (EOQ), the total cost per unit time
and the cycle time.
1
Expert's answer
2014-11-12T12:14:51-0500
D = 7200 kg per year,c = 20. K = 72, h = 25% of the investment in the inventories.
Economic order quantity is the order quantity that minimizes total inventory holding costs and ordering costs.
EOQ = (2D*K/h)^0.5 = (2*7200*72/0.25*20)^0.5 = 455.4
Total cost per unit = purchase cost or production cost + ordering cost + holding cost = 20 + 72 + 20*0.25 = 97
Cycle time = 311/(D/EOQ) = 311/(7200/455.4) = 19.7 days.